Solve by factoring: 15 x^2 + 11 x - 14 = 0
x =
Enter ONLY ONE of the 2 possible answers.
Write answer as an integer or reduced fraction.

Answer by stanbon(75887)

Solve by factoring: 15*x^2 + 11 x - 14 = 0
(3x-2)(5x+7) = 0
------
x = 2/3
---------------
Enter ONLY ONE of the 2 possible answers.
Write answer as an integer or reduced fraction.
===============
Cheers,
Stan H.
---------------